# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 3.7 Correctness
Proof 3.11.
The aggregated response is $\mathbf{z}=\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}=\sum_{i}\lambda_{i}(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})=\sum_{i}\lambda_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}=\mathbf{y}+c\mathbf{s}_{1}$. Both $\{\mathbf{y}_{i}\}$ and $\{\mathbf{s}_{1,i}\}$ are degree-(T1)(T\!-\!1) Shamir shares of $\mathbf{y}$ and $\mathbf{s}_{1}$ respectively, evaluated at the same points, so Lagrange reconstruction yields both correct secrets simultaneously. The masked commitment aggregation gives $\mathbf{W}=\sum_{i}\mathbf{W}_{i}=\sum_{i}(\lambda_{i}\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)})=\mathbf{A}\sum_{i}\lambda_{i}\mathbf{y}_{i}=\mathbf{A}\mathbf{y}=\mathbf{w}$, where mask cancellation follows from Lemma 3.4. Since $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ and $\mathbf{w}=\mathbf{A}\mathbf{y}$, verification proceeds as in single-signer ML-DSA.
@@ 3.8 Nonce Distribution and Security
The aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (the sum of each party’s constant-term contribution from the nonce DKG, equivalently $\mathbf{y}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ by Lagrange reconstruction) follows an Irwin-Hall distribution concentrated around zero. This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we compute per-coordinate $R^{\epsilon}_{2,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$. Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R^{\epsilon}_{2,\mathsf{coord}})^{1280}\approx 2^{5.4}$ for $|S|\leq 17$ ($\approx$3-bit security loss). For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss). See Theorem 4.11 and Appendix D.
@@ 3.10 The r0-Check Subprotocol
We address this via three deployment profiles (Table 2). In Profile P1 (TEE), $c\mathbf{s}_{2}$ is computed inside the enclave and never exposed. In Profile P2 (MPC), parties jointly evaluate the r0-check without reconstructing $c\mathbf{s}_{2}$. In Profile P3+ (2PC), two CPs each hold an additive share, so neither learns the complete value. See Section 6 for P2/P3 details.
@@ 4 Security Analysis
We prove that our masked threshold ML-DSA protocol achieves EUF-CMA security and privacy. This section contains complete proofs; we do not defer arguments to supplementary material.
@@ 4.1 Security Model
With Shamir nonce DKG and the coordinator-based communication model (Remark 3.7), Profiles P1 and P3+ require signing sets of size only $|S|\geq T$. The honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has high conditional min-entropy (Theorem 4.4) from the adversary’s view, vastly exceeding the secret key entropy. The key privacy argument rests on what the adversary observes:
@@
    Response $\mathbf{z}_{h}$: Statistical privacy via nonce DKG (Theorem 4.4). Requires no computational assumptions, holds even with $|S\setminus C|=1$.
@@
If $\mathbf{W}_{h}$ were broadcast with $|S\setminus C|=1$, the pairwise mask $\mathbf{m}_{h}^{(w)}$ would be computable by the adversary (all seeds involve corrupted counterparties), exposing $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$. Since $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k>\ell$ is generically injective (under the NTT isomorphism, a random $k\times\ell$ matrix over $\mathbb{Z}_{q}$ has full column rank with probability $\geq 1-\ell/q$), the adversary could recover $\mathbf{y}_{h}$ via linear algebra—this is not a Module-LWE instance because $\mathbf{y}_{h}$ has full-size (not short) coefficients. Combined with $\mathbf{z}_{h}$ (extractable from the public signature), this would yield $\mathbf{s}_{1,h}=c^{-1}(\mathbf{z}_{h}-\mathbf{y}_{h})$, a complete key recovery. The coordinator-only communication model eliminates this attack vector entirely.
@@ 4.2 Unforgeability
If Module-SIS is hard and $H$ is a random oracle, then the masked threshold ML-DSA scheme is EUF-CMA secure:
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
where $q_{H},q_{s}$ are random oracle and signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence with per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (Theorem 4.11). The $\sqrt{R_{2}^{\mathsf{vec}}}$ factor arises from the game hop between Irwin-Hall and uniform nonce distributions via Cauchy-Schwarz; for $|S|\leq 17$, $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$, giving $\sqrt{R_{2}^{\mathsf{vec}}}\approx 2^{2.7}$ ($\approx$3-bit loss). For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss). The loss matches the per-coordinate composition inherent to the generic Rényi proof technique [Raccoon2024].
Proof 4.3.
@@
The real EUF-CMA game. Challenger runs $\mathsf{KeyGen}$ honestly, gives $\mathcal{A}$ the public key $\mathsf{pk}$ and corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$.
@@
During signing, sample challenge $c\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{C}$ uniformly before computing $\mathbf{w}_{1}$, then program $H(\mu\|\mathbf{w}_{1}):=c$. This fails only on collision. By birthday bound: $|\Pr[\mathsf{Win}_{2}]-\Pr[\mathsf{Win}_{1}]|\leq(q_{H}+q_{s})^{2}/2^{256}$.
@@
Replace honest parties’ shares with uniformly random values, setting $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $j\notin C$. By $(T,N)$-Shamir security (any <T<T shares reveal nothing): $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
@@
Replace the Irwin-Hall nonce distribution with uniform nonces: in each signing session, sample $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform instead of $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (Irwin-Hall). By the smooth Rényi divergence bound (Theorem 4.11 with $\alpha=2$):
$\Pr[\mathsf{Win}_{3}]\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
where $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, $n\ell=1280$), giving $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ for $|S|\leq 17$. The tail probability satisfies $\epsilon_{\mathsf{IH}}<10^{-30}$. This is the only game hop that incurs the Irwin-Hall security loss; individual shares $\mathbf{z}_{j}$ have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
@@
Starting from uniform nonces (Game 3.5), simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$, and program $H(\mu\|\mathbf{w}_{1}):=c$. The masked values $\mathbf{W}_{j}$ are simulated as uniform by Lemma 4.9. By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.

### v5

@@ 3.7 Correctness
Proof 3.13.
The aggregated response is $\mathbf{z}=\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}=\sum_{i}\lambda_{i}(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})=\sum_{i}\lambda_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}=\mathbf{y}+c\mathbf{s}_{1}$. Both $\{\mathbf{y}_{i}\}$ and $\{\mathbf{s}_{1,i}\}$ are degree-(T1)(T\!-\!1) Shamir shares of $\mathbf{y}$ and $\mathbf{s}_{1}$ respectively, evaluated at the same set of points: KeyGen assigns $\mathbf{s}_{1,j}=p_{\mathbf{s}_{1}}(j)$ using party indices $j\in[N]$, and the nonce DKG assigns $\mathbf{y}_{j}=\mathbf{F}(j)$ using the same evaluation domain (Algorithm 2). Hence the Lagrange coefficients $\lambda_{i}$ for the signing set $S$ apply equally to both sharings, and reconstruction yields both correct secrets simultaneously. The masked commitment aggregation gives $\mathbf{W}=\sum_{i}\mathbf{W}_{i}=\sum_{i}(\lambda_{i}\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)})=\mathbf{A}\sum_{i}\lambda_{i}\mathbf{y}_{i}=\mathbf{A}\mathbf{y}=\mathbf{w}$, where mask cancellation follows from Lemma 3.6. Since $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ and $\mathbf{w}=\mathbf{A}\mathbf{y}$, verification proceeds as in single-signer ML-DSA.
@@ 3.8 Nonce Distribution and Security
The aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (the sum of each party’s constant-term contribution from the nonce DKG) follows an Irwin-Hall distribution concentrated around zero. This equals the Lagrange reconstruction $\mathbf{y}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ by the following two-step argument: letting $\mathbf{F}(x)=\sum_{k\in S}\mathbf{f}_{k}(x)$ be the combined degree-(T1)(T\!-\!1) polynomial, we have $\mathbf{F}(0)=\sum_{k\in S}\mathbf{f}_{k}(0)=\sum_{k\in S}\hat{\mathbf{y}}_{k}$ (direct, by definition of the constant term), and also $\mathbf{F}(0)=\sum_{i\in S}\lambda_{i}\mathbf{F}(i)=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ (Lagrange interpolation at 0). This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we compute per-coordinate $R^{\epsilon}_{2,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$. Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R^{\epsilon}_{2,\mathsf{coord}})^{1280}\leq(1.003)^{1280}\approx 2^{5.5}$ for $|S|\leq 17$ ($\approx$3-bit security loss; exact: $(1.00292)^{1280}\approx 2^{5.4}$). For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss). See Theorem 4.17 and Appendix D.
@@ 3.10 The r0-Check Subprotocol
We address this via three deployment profiles (Table 2). In Profile P1 (TEE), $c\mathbf{s}_{2}$ is computed inside the enclave and never exposed; $c\mathbf{t}_{0}$ is derived from the formula above, and the hint $\mathbf{h}$ is computed inside the enclave before output. In Profile P2 (MPC), parties jointly evaluate the r0-check without reconstructing $c\mathbf{s}_{2}$: the MPC circuit extends to also compute $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+c\mathbf{s}_{2}-c\mathbf{t}_{1}\cdot 2^{d}$ using the internally-held intermediate $c\mathbf{s}_{2}$, and outputs $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\,\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$ directly (since $\mathsf{MakeHint}$ is a bitwise comparison of bounded-norm vectors, it adds only $O(knd)$ comparison gates to the r0-check circuit); $c\mathbf{s}_{2}$ is never exposed outside the MPC. In Profile P3+ (2PC), two CPs each hold an additive share of $c\mathbf{s}_{2}$, so neither learns the complete value; hint computation proceeds analogously within the 2PC. See Section 6 for P2/P3 details.
@@ 4 Security Analysis
We prove that our masked threshold ML-DSA protocol achieves EUF-CMA security and statistical share privacy. This section contains proofs for all main results; full UC proofs for P2 and P3+ appear in Appendix H.
@@ 4.1 Security Model
With Shamir nonce DKG, Profiles P1 and P3+ require only $|S|\geq T$. The adversary observes: (1) nonce shares $\mathbf{y}_{h}$ with high conditional min-entropy (statistical, Theorem 4.7); (2) r0-check shares $\mathbf{V}_{i}$ never broadcast, staying inside the TEE/2PC; (3) masked commitments $\mathbf{W}_{i}$ sent only to the coordinator, which publishes only $\mathbf{w}_{1}$ (already implicit in $\sigma$). Individual $\mathbf{W}_{i}$ must stay private: if broadcast with $|S\setminus C|=1$, the adversary could recover $\mathbf{y}_{h}$ from $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$ via linear algebra (since $\mathbf{A}$ is generically injective [FIPS204, LPR10]), then extract $\mathbf{s}_{1,h}$. In P2, mask hiding (Lemma 4.15) protects broadcast $\mathbf{W}_{i}$ under the existing $|S\setminus C|\geq 2$ condition; note that mask hiding applies only to $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ (which do not appear in the signature) and does not contribute to nonce share privacy—the latter is provided purely by the statistical properties of the nonce DKG (Theorem 4.7). Additionally, P2’s blame protocol (Section 6.1) assumes challenge invertibility ($c$ invertible in $R_{q}$; empirically $99.9\%$ of ML-DSA-65 challenges are invertible [FIPS204]); this invertibility claim is heuristic and does not constitute a formal negligibility bound.
@@
    Nonce share $\mathbf{y}_{h}$: Statistical privacy via nonce DKG (Theorem 4.7). The nonce share has high conditional min-entropy (no computational assumptions), holding even with $|S\setminus C|=1$. Key privacy given the signature $\sigma$ is computational and not formally reduced to Module-LWE in this work; it remains an open problem to provide a tight security reduction for the bounded-distance decoding problem $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ when $c$ is a sparse challenge and $\mathbf{y}_{h}$ has bounded (but not short) coefficients. We note that single-signer ML-DSA faces the identical question: FIPS 204 [FIPS204] does not formally reduce key privacy (given signatures) to Module-LWE either.
@@
If $\mathbf{W}_{h}$ were broadcast with $|S\setminus C|=1$, the pairwise mask $\mathbf{m}_{h}^{(w)}$ would be computable by the adversary (all seeds involve corrupted counterparties), exposing $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$. Since $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k>\ell$ is generically injective (the NTT isomorphism $R_{q}\cong\mathbb{Z}_{q}^{n}$ [FIPS204] decomposes $\mathbf{A}$ into $n$ independent $k\times\ell$ matrices over $\mathbb{Z}_{q}$; we heuristically model each as uniformly random and apply the $\mathbb{Z}_{q}$ full-rank bound of [LPR10, Lemma 2.1] componentwise, giving full column rank with probability $\geq 1-q^{\ell-k}$; this is the same standard heuristic used in ML-DSA [DKL18, FIPS204]), the adversary could recover $\mathbf{y}_{h}$ via linear algebra—this is not a Module-LWE instance because $\mathbf{y}_{h}$ has full-size (not short) coefficients (i.e., coefficients drawn from $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]$ with $\lfloor\gamma_{1}/|S|\rfloor\approx 2^{19}/|S|\gg\eta=4$; the Module-LWE hardness assumption requires error terms with small coefficients). Combined with $\mathbf{z}_{h}$ (extractable from the public signature), this would yield $\mathbf{s}_{1,h}=c^{-1}(\mathbf{z}_{h}-\mathbf{y}_{h})$, a complete key recovery (assuming $c$ is invertible in $R_{q}$; empirically, $99.9\%$ of ML-DSA-65 challenges are invertible [FIPS204]). The coordinator-only communication model eliminates this attack vector entirely. Caveat for Profile P2 blame path: In P2, the distributed blame protocol (Section 6.1) also relies on $c$-invertibility to verify party responses without a coordinator. The claim that $99.9\%$ of challenges are invertible is an empirical estimate and does not yield a cryptographically negligible failure probability; this invertibility argument for P2 blame verification is therefore heuristic, not a formal negligibility bound.
@@ 4.2 Unforgeability
If Module-SIS is hard and $H$ is a random oracle, then the masked threshold ML-DSA scheme is EUF-CMA secure. Using the direct shift-invariance bound (Corollary Q.24, Theorem Q.25), the Irwin-Hall nonce distribution introduces at most $q_{s}\times 6.1\times 10^{-3}$ bits of security loss for $|S|\leq 17$ ($q_{s}\times 0.013$ bits for $|S|\leq 33$), where $q_{s}$ is the number of signing queries. For $q_{s}\leq 1{,}600$ queries with $|S|\leq 17$, this is $\leq 9.8$ bits, yielding a proven security bound of ${\approx}86$ bits under MSIS (=969.8=96-9.8; the 96-bit baseline is a proof-technique artifact from Cauchy-Schwarz halving the 192-bit NIST Level 3 MSIS hardness once, identical to single-signer ML-DSA). The following bound holds via Rényi divergence:
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})$ $\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}}$
$\quad+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
where $q_{H},q_{s}$ are random oracle and signing queries, $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{n\ell}$ is the full-vector per-session Rényi divergence from the direct shift-invariance analysis (here $n=256$ is the ring degree, $\ell=5$, giving $n\ell=1280$) with per-coordinate $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ for $|S|=17$ (Corollary Q.24), giving $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (smooth Rényi tail parameter, Theorem 4.17; note the per-coordinate empirical tail probability from the FFT computation is $2.3\times 10^{-20}$ for $|S|=17$, Corollary Q.24—these are distinct quantities: the former is the smooth Rényi parameter used in the formal framework, the latter is the direct tail probability). The $(R_{2}^{\mathsf{vec,shift}})^{q_{s}/2}$ factor accounts for multi-session Rényi composition (Remark Q.17): each of the $q_{s}$ signing sessions contributes one factor of $R_{2}^{\mathsf{vec,shift}}$ to the joint divergence. The total security loss is $\approx 6.1\times 10^{-3}\cdot q_{s}$ bits, giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (non-vacuous for $q_{s}<16{,}000$; the 96-bit figure assumes $q_{H}=1$—see Remark 4.6 for the general $q_{H}$-dependent form, which is identical to the single-signer FIPS 204 convention). Our threshold construction adds only $\leq 10$ bits of additional EUF-CMA loss from the Irwin-Hall nonce distribution for $q_{s}\leq 1{,}600$. The conservative Raccoon-style analytical bound is given in Remark 4.5.
Corollary 4.3 (Irwin-Hall Nonce Security Loss).
Under the conditions of Theorem 4.2, using the direct shift-invariance bound (Theorem Q.25, Corollary Q.24), the EUF-CMA security loss from Irwin-Hall nonces is $0.007\cdot q_{s}$ bits for $|S|\leq 17$ and $0.013\cdot q_{s}$ bits for $|S|\leq 33$, where $q_{s}$ is the number of signing queries (multi-session composition; Remark Q.17). For $q_{s}=1$ (single session), the per-session loss is <0.007<0.007 bits ($|S|\leq 17$) or <0.013<0.013 bits ($|S|\leq 33$). The single-session bound first exceeds $1$ bit only at $|S|=2584$, fully eliminating scalability as a per-session security concern.
Proof 4.4.
@@
The real EUF-CMA game. Challenger runs $\mathsf{KeyGen}$ honestly, gives $\mathcal{A}$ the public key $\mathsf{pk}$ and corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$. Signing oracle queries may return $\bot$ (rejection abort); $\mathcal{A}$ may re-issue the same or different queries. Security bounds are in terms of query counts $q_{s}$; each attempt succeeds with non-negligible probability $\epsilon_{\mathsf{acc}}\in[21\%,45\%]$ depending on $|S|$ (see Table 5; comparable to single-signer ML-DSA’s $\approx 25\%$), so all $q_{s}$ successful queries complete in expected $O(q_{s}/\epsilon_{\mathsf{acc}})=O(q_{s})$ attempts.
@@
During signing, sample challenge $c\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{C}$ uniformly before computing $\mathbf{w}_{1}$, then program $H(\mu\|\mathbf{w}_{1}):=c$. This fails only on collision. By birthday bound: $|\Pr[\mathsf{Win}_{2}]-\Pr[\mathsf{Win}_{1}]|\leq(q_{H}+q_{s})^{2}/2^{256}$, where $2^{256}$ is the hash output space size (SHAKE-256 gives 256-bit challenges; this is a conservative bound since $|C|<2^{256}$).
@@
The adversary’s view reveals no information about honest parties’ key shares: by $(T,N)$-Shamir perfect hiding, the joint distribution of the T1T-1 corrupted evaluations is statistically independent of the secret $\mathbf{s}_{1}$ (though the evaluations themselves are correlated as points on a shared polynomial).
Important clarification: The signing oracle does not change in this hop — it continues using the actual short key shares $\{\mathbf{s}_{1,j},\mathbf{s}_{2,j}\}_{j\notin C}$ internally. Replacing those shares with $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ would cause the z-bound check ($\|\mathbf{z}\|_{\infty}\geq\gamma_{1}-\beta$ almost surely for large $\mathbf{s}_{1,j}$) and the r0-check to fail with overwhelming probability, breaking correctness. This argument is purely about what the adversary observes: because honest shares are never directly visible to the adversary (see profile-specific paragraph below), the adversary’s view is limited to the $|C|\leq T-1$ corrupted shares $\{\mathbf{s}_{1,i}\}_{i\in C}$ and $\{\mathbf{s}_{2,i}\}_{i\in C}$. By Shamir’s perfect hiding, T1T-1 evaluations of a degree-(T1)(T-1) polynomial reveal no information about the secret $\mathbf{s}_{1}=\mathbf{f}(0)$: the conditional distribution of $\mathbf{f}(0)$ given any T1T-1 evaluations is identical to the prior (marginal) distribution of $\mathbf{f}(0)$ (statistical distance zero: no event in the adversary’s view distinguishes Game 2 from Game 3). A simulator in Game 3 samples corrupted shares $\{\mathbf{s}_{1,i}\}_{i\in C}$ from the conditional Shamir distribution given $\mathbf{s}_{1}$ (computed via Lagrange interpolation with T|C|T-|C| uniformly random high-degree coefficients), which is identically distributed to the real key generation. Hence the adversary’s view in $\mathsf{Game}_{3}$ is identically distributed to that in $\mathsf{Game}_{2}$, giving $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
Profile-specific observability of $\mathbf{z}_{j}$. In Profiles P1 and P3+, individual signer responses $\mathbf{z}_{j}$ are transmitted only to the TEE coordinator or the honest CP, respectively, and are never observed by the adversary; the game hop is trivial. In Profile P2, $\mathbf{z}_{j}$ is contributed to $\mathcal{F}_{\mathsf{SPDZ}}$ for aggregation, which provides input privacy: only the aggregate $\mathbf{z}=\sum_{j\in S}\lambda_{j}\mathbf{z}_{j}$ is revealed. Since $\mathcal{F}_{\mathsf{SPDZ}}$ hides individual $\mathbf{z}_{j}$, the adversary gains no information about $\mathbf{s}_{1,j}$ through $\mathbf{z}_{j}$ in any profile. The Shamir replacement is therefore valid across all deployment profiles.
@@
Replace the Irwin-Hall nonce distribution with uniform nonces: in each signing session, sample $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform instead of $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (Irwin-Hall). The $q_{s}$ signing sessions use independent nonce DKGs, so the per-session nonce distributions are i.i.d. copies of the Irwin-Hall distribution (convolution of $|S|$ uniform shares). By the product rule for Rényi-2 divergence (where $R_{2}(P\|Q)=\sum_{x}P(x)^{2}/Q(x)$ in the ratio-sum convention): for independent sessions,
$R_{2}^{\mathsf{joint}}\;=\;\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}},$
since the joint divergence factors as a product of per-session divergences (see Remark Q.17 for explicit derivation). By the Rényi-2 security bound (Theorem 4.17 with $\alpha=2$, applied to the $q_{s}$-session joint distribution):
$\Pr[\mathsf{Win}_{3}]\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
where $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{n\ell}$ is the full-vector Rényi divergence per session from the direct shift-invariance analysis (Theorem Q.25), with per-coordinate $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ for $|S|=17$ (Corollary Q.24, $n\ell=1280$), giving $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$ for $|S|=17$. The tail probability satisfies $\epsilon_{\mathsf{IH}}<10^{-30}$ for $|S|\leq 17$ (this bound is $|S|$-dependent; see Theorem 4.17 for the general formula as a function of $|S|$, $n$, and $\gamma_{1}$). This is the only game hop that incurs the Irwin-Hall security loss; nonce shares $\mathbf{y}_{j}$ have high min-entropy by Theorem 4.7 (statistical, no computational assumptions). Note on $\epsilon_{\mathsf{IH}}$: The value $\epsilon_{\mathsf{IH}}<10^{-30}$ above is the smooth Rényi tail parameter from Raccoon Lemma 4.2 (Theorem 4.17, formal analytical bound). The $\epsilon$ column in Corollary Q.24 ($2.3\times 10^{-20}$ for $|S|=17$) is the per-coordinate empirical tail probability $\Pr[|Y_{\mathsf{coord}}|>\gamma_{1}-\beta]$ from the Direct Shift FFT computation—a distinct, looser quantity; both are negligible for all practical $q_{s}$. See Remark Q.17 for the concrete multi-session security accounting. The conservative Raccoon-style alternative (Remark 4.5) gives $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ via the analytical bound, which is vacuous at $q_{s}\gtrsim 36$.
@@
Starting from uniform nonces (Game 3.5), simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$, set $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{r},2\gamma_{2})$, program $H(\mu\|\mathbf{w}_{1}):=c$, and set hint $\mathbf{h}=\mathbf{0}$ (zero hint; since $\mathsf{UseHint}(\mathbf{0},\mathbf{r},2\gamma_{2})=\mathsf{HighBits}(\mathbf{r},2\gamma_{2})=\mathbf{w}_{1}$, verification passes with hint-weight $\|\mathbf{h}\|_{1}=0\leq\omega$).
Remark on the hint distribution. The simulated hint $\mathbf{h}=\mathbf{0}$ differs in distribution from the real hint $\mathbf{h}_{\mathsf{real}}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\,\mathbf{r},\,2\gamma_{2})$ in Game 3.5, which may carry up to $\omega=55$ nonzero bits. This substitution does not increase the adversary’s forging advantage. In Game 3.5 the real hint encodes partial information about the secret key component $\mathbf{t}_{0}$ (specifically, the “carry” pattern of $\mathsf{LowBits}(\mathbf{r},2\gamma_{2})$ relative to $\pm\gamma_{2}$, which depends on $c\mathbf{t}_{0}-c\mathbf{s}_{2}$); by contrast, $\mathbf{h}=\mathbf{0}$ reveals nothing about the key. The adversary’s oracle view in Game 4 is therefore strictly less informative about the secret key than in Game 3.5, giving $\Pr[\mathsf{Win}_{4}]\leq\Pr[\mathsf{Win}_{3.5}]$ with no additional loss from the hint distribution change. (This is a standard observation in Dilithium-family proofs [FIPS204]: replacing a real hint with $\mathbf{0}$ is a valid simulation that removes rather than adds key leakage.)
The masked values $\mathbf{W}_{j}$ require separate treatment by profile. In P1 and P3+, individual $\mathbf{W}_{j}$ are transmitted only to the TEE coordinator or CP respectively and are never observable by the adversary; the adversary’s view contains only the aggregate $\mathbf{w}_{1}=\mathsf{HighBits}(\sum_{j}\mathbf{W}_{j},2\gamma_{2})$, which is already implicit in $\sigma$ — so the simulator need not produce individual $\mathbf{W}_{j}$ at all, and Lemma 4.15 is not invoked for these profiles. In P2, parties broadcast individual $\mathbf{W}_{j}$; the simulator produces each as uniform by Lemma 4.15, which applies since $|S\setminus C|\geq 2$ is a stated requirement of P2. By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (there are $\binom{N}{2}$ pairwise PRF seeds established at key generation; a union bound over all honest–honest seed pairs gives this term, independent of $q_{s}$ since each domain-separated session ID yields a fresh PRF input; for P1 and P3+ this term is vacuously 0 as the adversary does not observe individual $\mathbf{W}_{j}$, so the bound is conservative for those profiles). The change in $\mathbf{z}$-distribution between Game 3.5 (truncated shifted uniform: $\mathbf{y}+c\mathbf{s}_{1}$ conditioned on $\|\mathbf{y}+c\mathbf{s}_{1}\|_{\infty}<\gamma_{1}-\beta$) and Game 4 (direct uniform on $\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$) is absorbed by the SelfTargetMSIS reduction [KLS18]: the adversary $\mathcal{B}$ extracts a valid SIS solution from any forgery $(\mathbf{z}^{*},c^{*},\mathbf{h}^{*})$ satisfying $\|\mathbf{z}^{*}\|_{\infty}<\gamma_{1}-\beta$, and this norm check is enforced identically in both games regardless of which simulated distribution $\mathbf{z}$ was drawn from.